Indirect measurement: a building is 22ft high and casts a 40ft shadow. Mr.berg casts a 10ft shadow. The triangle formed by the building and it's shadow is similar to the triangle formed by Mr.berg and his shadow. How tall is Mr.berg?

Answer by Cromlix(4381)

Height of building = 22ft
Length of shadow = 40ft
Height of Mr. Berg = ?
Length of shadow = 10ft
10/40 = ?/22
Cross multiply
220 = 40?
Height of Mr. Berg = 220/40 = 5.5ft
or 5ft 6 inches
Hope this helps.
:-)